OpenStudy (anonymous):

it is system of equation if x+2y=x+4 and 9y=-3x+9+9y what is the answer cant you explain it please??????????????????? :) ;)

OpenStudy (anonymous):

For the first one, subtract x from both sides. For the second one, subtract 9y from both sides.

OpenStudy (anonymous):

I am pretty sure there is a typo somewhere in the equations. Can you please recheck?

OpenStudy (anonymous):

Yes it did look a little strange..

OpenStudy (anonymous):

okay let me check that is the question that the teacher give me

OpenStudy (openstudier):

Hint: Try to get both equations into this form first: \[ax+by=c\]

OpenStudy (anonymous):

i think u should first get rid of the x in da first equation then ull juz solve for y then plug it in for da next equation nd solve for x

OpenStudy (anonymous):

x+2y=x+4 if then 2y =4 and hence y=4/2=2 9y=-3x+9+9y 9y -9y=-3x+9+9y-9y or 0=-3x+9 or 3x=9 or x=9/3=3